# MN-TSG: Continuous Time Series Generation with Irregular Observations

### v1

@@ Abstract
Time series generation (TSG) plays a critical role in a wide range of domains, such as healthcare. However, most existing methods assume regularly sampled observations and fixed output resolutions, which are often misaligned with real-world scenarios where data are irregularly sampled and sparsely observed. This mismatch is particularly problematic in applications such as clinical monitoring, where irregular measurements must support downstream tasks requiring continuous and high-resolution time series. Neural Controlled Differential Equations (NCDEs) have shown strong potential for modeling irregular time series, yet they still face challenges in capturing complex dynamic temporal patterns and supporting continuous TSG. To address these limitations, we propose MN-TSG, a novel framework that explores Mixture-of-Experts (MoE)–based NCDEs and integrates them with existing TSG models for irregular and continuous generation tasks. The core of MN-TSG lies in a MoE-NCDE architecture with dynamically parameterized expert functions and a decoupled design that facilitates more effective optimization of MoE dynamics. Furthermore, we leverage existing TSG models to learn the joint distribution over the mixture of experts and the generated time series. This enables the framework not only to generate new samples, but also to produce appropriate expert configurations tailored to each sample, thereby supporting refined continuous TSG. Extensive experiments on ten public and synthetic datasets demonstrate the effectiveness of MN-TSG, consistently outperforming strong TSG baselines on both irregular-to-regular and irregular-to-continuous generation tasks. The code is available at the link https://github.com/microsoft/TimeCraft/tree/main/MNTSG.
Keywords Irregular time series continuous time series generation deep learning architecture
@@ 1 Introduction
Time series generation (TSG) plays a pivotal role across a diverse range of real-world domains, such as healthcare (Deng et al., 2025), finance (Vasiliu et al., 2024), energy systems (Fuest et al., 2025), and climate science (Ye et al., 2024) by producing synthetic sequences that facilitate data augmentation for low-resource contexts (Tang et al., 2025), enable privacy‑preserving data sharing (Tian et al., 2024), and support high‑fidelity simulations for critical decision‑making (Turowski et al., 2024). While numerous TSG methods have been developed, they generally build upon a fundamental assumption that the time series (both training data and generated data) are regular, which is represented as discrete observations with equally spaced timestamps. As depicted in Fig. 1(a), most existing works treat training sequences as uniformly observed and produce outputs sampled at fixed intervals (Mushunje et al., 2023; Li, 2023; Huang et al., 2025; Ge et al., 2025). As shown in Fig. 1(b), few approaches relax the first assumption—accepting irregularly observed training data—but still generate outputs with fixed intervals (Nikitin et al., 2024; Ramponi et al., 2018). This mismatch contrasts sharply with real-world scenarios, where data often arrives irregularly, yet downstream applications demand continuous, high-resolution sequences. For example, in clinical settings like ICU monitoring, electronic health record entries may be recorded at irregular and sparse intervals, sometimes with up to 80% missingness in some features, but tasks like sepsis forecasting or patient trajectory simulation require continuous, temporally dense data (Tian et al., 2024; Li et al., 2025). Thus, continuous time series generation from irregular observations is a pressing and under-addressed challenge within the TSG community.
@@
Continuous time series modeling and the challenge of irregular observations have been actively studied, with several techniques proposed. Among all the techniques, Neural Ordinary Differential Equations (NODE) and its following work, Neural Controlled Differential Equations (NCDE) have been particularly popular, which we will refer to as NCDE-based methods. They combine differential equations in continuous time with learnable neural networks to approximate temporal dynamics. This makes NCDE-based methods especially suitable for continuous TS modeling and handling irregular observations. Some recent studies have leveraged NCDE-based methods for TSG (Lim et al., 2023; Naiman et al., 2024), typically following a two-stage paradigm: they first transform irregularly sampled TS into regularly sampled ones using NCDEs or their variants, and then apply standard TSG models such as GANs (Jeon et al., 2022) and VAEs (Naiman et al., 2024) to generate data on a regularly sampled TS.
However, on the one hand, they only focus on using NCDE to process irregular inputs and then generate regular time series (Fig. 1(b)), without in-depth research on how to generate and evaluate continuous time series (Fig. 1(c)), which still needs further exploration. On the other hand, existing methods simply apply the standard NCDE without making any improvements. However, the performance of the standard NCDE can still be further improved, which is helpful for dealing with irregular data and continuous generation tasks. We have noticed that the standard NCDE faces challenges in learning dynamically changing temporal patterns, especially in the generation of time series with different classes. Moreover, the standard NCDE involves the optimization of multiple components, and end-to-end training may cause the model to fail to focus on the learning of dynamic functions.
In this study, we propose a novel TSG framework, MN-TSG, which explores Mixture of Experts Neural Controlled Differential Equations (MoE-NCDE) and integrates it with existing TSG models for continuous time series generation. To address the aforementioned challenges, we introduce three key architectural innovations. First, we propose a Mixture of Experts (MoE) based dynamic functions for NCDE to effectively capture the diverse dynamics inherent in real-world time series data. Second, we decouple the model architecture to separately learn time series dynamics and the mapping between raw observations and hidden states by pretraining a channel-wise autoencoder. Third, we propose a joint learning approach that simultaneously models discrete observations and underlying MOE dynamics within the TSG framework. Our contributions are as follows:
@@
We propose the decoupled MOE Neural CDE, which improves the standard NCDE by proposing the mixture-of-experts dynamic functions (MOE dynamics) and a decoupled design to help NCDE better focus on the learning of MOE dynamics and handling dynamically changing patterns.
We propose the continuous time series generation framework called MN-TSG. It learns the joint distribution of dynamic functions and time series samples based on existing generative models. This allows the model to generate new samples while assigning them suitable dynamic function weights, leading to more reasonable continuous time series generation.
We conduct comprehensive experiments with advanced baselines on four popular TSF datasets, four medical ECG datasets, and two synthetic datasets. In terms of quantitative and qualitative evaluations, results show that our design enhances NCDE’s ability to learn dynamically changing patterns from irregular temporal data, achieving better performance in both irregular-to-regular and irregular-to-continuous generation tasks.
@@ 2 Related Work
The most relevant methods are NCDE-based generators, e.g., KOVAE (Jeon et al., 2022) and GTGAN (Naiman et al., 2024), which build on standard NCDE and conduct irregular-to-regular generations. Some approaches also attempt to use attention-style mechanisms to capture the relationship between the observed sparse values and the unknown target time-steps, learning the time dependencies and thereby enabling prediction at arbitrary time-steps (Shukla and Marlin, ; Yalavarthi et al., 2025). Unlike them, we improve the standard NCDE and propose the decoupled MOE NCDE. Moreover, we design a framework for continuous time series generation and explore its evaluations. Complete related work about TSG models for regular and irregular data is provided in the Appendix A.1.
@@ 3 Method
Overview. Figure 2 illustrates MN-TSG. Training: MOE-NCDE is trained on irregular samples using mixture-of-experts dynamic functions to better capture dynamic patterns. Further, a pretrained channel-wise autoencoder replaces and the reconstruction function with a frozen channel-wise encoder (CE) and frozen channel-wise decoder (CD) to make MOE-NCDE focus on learning MOE dynamics. After training, MOE-NCDE yields imputed regular samples and MOE weights of different dynamic functions, which are then fed into a diffusion model to join training and jointly generate new samples with corresponding expert weights. Generation: The generated samples and expert weights are fed back into MOE-NCDE for continuous generation, producing refined, longer time series. Finally, the right half compares traditional NCDE with our decoupled MOE-NCDE.
@@
Notation. Suppose we have observed multivariate time series. Each time series , for , is a continuous multivariate function defined over its time interval, where is the number of channels/variables. In practice, we observe each multivariate time series at discrete and irregular time points , where . The observed dataset is:
@@
Problem Definition. Given dataset 111We assume the underlying dynamic of the data is continuous and differentiable. Some discussion is made in Section 5., the goal is to learn a generative model over discrete and irregular time series for irregular or continuous generation, i.e., our goal is to learn a model capable of sampling new sequences . The generated dataset is denoted as where denotes arbitrary time step of -th sample.
3.1 Neural Controlled Differential Equations
Neural CDEs (Kidger et al., 2020) generalize Neural ODEs (Chen et al., 2018) by modeling the latent dynamics as being driven by a continuous input path (control path). The CDE is defined as:
@@ 3.1 Neural Controlled Differential Equations
where is a learnable neural dynamics function, and is a continuous interpolation (e.g., a natural cubic spline) of the discrete observations . is the initial latent state mapped from based on a learnable neural network. represents the hidden state at time step , and a reconstruction function subsequently maps this hidden representation back to the original channel space to produce the predicted time-series values.
This interpolation defines the control path that drives the evolution of the latent state . CDESolver is used to solve controlled differential equations, which can be formalized as:
3.2 Modeling Diverse Dynamics with Mixture of Experts
Real-world time series data often exhibit diverse dynamics across different samples, making it difficult for a single model to capture all behaviors effectively. By leveraging a Mixture of Experts (MoE) framework, we allow the model to assign specialized dynamics to different samples, enabling more expressive and personalized temporal modeling. This flexibility enhances both the representational capacity and the generalization of the learned dynamics.
@@ 3.2 Modeling Diverse Dynamics with Mixture of Experts
The MoE dynamics are composed of a router and dynamic functions (each is parameterized as a neural network), where . The router takes (potentially irregular) observations as input and outputs a routing score for each expert. Specifically, we first use the cubic interpolation function to impute the irregular time series and then feed them into the router as observations to obtain all experts’ scores for the -th sample. We implement the router by using simple MLPs. Then, a softmax function is used over the scores to compute a dense routing distribution. Specifically, given for all experts’ scores of -th sample, the final dynamics function is then computed as a weighted combination of all experts’ functions:
Finally, by substituting into standard NCDE formulas (Eq. 2 and Eq. 3), The Mixture-of-Experts-based MOE-NCDE is obtained.
3.3 Decoupled Training of MoE Dynamics
MOE-NCDE involves learning time series dynamic functions with a mixture of experts (MOE dynamics) and the mapping between raw observations and hidden states ( and reconstruction function). Existing methods often train all of them end-to-end, which can hinder MOE-dynamics training. Firstly, the quality of and the reconstruction function impact dynamic function learning. Secondly, jointly optimizing them prevents the model from focusing solely on the MOE dynamics.
To address MOE-NCDE training challenges, we propose a decoupled strategy: first pre-train and the reconstruction function using a channel-wise autoencoder with two lightweight MLPs. The training pipeline is shown in Algorithm 4. Specifically, for each batch, we keep the channel dimension, flatten the rest, and train the encoder to obtain hidden states and the decoder to reconstruct the channel dimension. Then, MOE-NCDE is trained in the condition of and the reconstruction function are replaced by the frozen Channel-wise Encoder (CE) and Channel-wise Decoder (CD). This way allows the model to focus solely on the learning of MOE-dynamics. The MoE dynamics can be trained through MSE loss based on CDESolver and the pretrained channel-wise autoencoder in the algorithm 1.
3.4 Joint Diffusion Modeling of Observations and Expert Dynamics
While prior diffusion-based models focus on learning the distribution of discrete time series observations, i.e., , they often overlook the underlying dynamics. By extending the diffusion model to jointly model both the observations and the corresponding MoE weights , we enable the diffusion model to generate not only observations but also appropriate expert weights for modeling temporal dynamics.
Notably, the diffusion model prefers data with a fixed length while the observed dataset might be irregular with changeable length. We leverage trained to impute the irregular data to a fixed length firstly. Then, the generation model that learn the joint distribution can be trained by the Appendix algorithm 2. In this way, we can obtain new samples and new expert weights . Finally, we substitute them into the CDE Solver, and for each generated sample, we perform continuous time series generation based on the respective generated expert weights. We show the process in the Appendix algorithm 3.
@@ 4 Experiments
We conduct extensive experiments using multiple real-world and synthetic datasets. The experiment goal is to investigate on the following research questions: Q1: How effective is our method in handling missing data and generating regular data? (Section 4.2) Q2: Can our method better perform the task of continuous time series generation? (Section 4.3) Q3: Are the relevant designs of our MOE-NCDE truly effective (e.g., MOE dynamic functions and decoupled design)? (Section 4.4)

### v2

# Diff-MN: Diffusion Parameterized MoE-NCDE for Continuous Time Series Generation with Irregular Observations

@@ Abstract
Time series generation (TSG) is widely used across domains, yet most existing methods assume regular sampling and fixed output resolutions. These assumptions are often violated in practice, where observations are irregular and sparse, while downstream applications require continuous and high-resolution TS. Although Neural Controlled Differential Equation (NCDE) is promising for modeling irregular TS, it is constrained by a single dynamics function, tightly coupled optimization, and limited ability to adapt learned dynamics to newly generated samples from the generative model. We propose Diff-MN, a continuous TSG framework that enhances NCDE with a Mixture-of-Experts (MoE) dynamics function and a decoupled architectural design for dynamics-focused training. To further enable NCDE to generalize to newly generated samples, Diff-MN employs a diffusion model to parameterize the NCDE temporal dynamics parameters (MoE weights), i.e., jointly learn the distribution of TS data and MoE weights. This design allows sample-specific NCDE parameters to be generated for continuous TS generation. Experiments on ten public and synthetic datasets demonstrate that Diff-MN consistently outperforms strong baselines on both irregular-to-regular and irregular-to-continuous TSG tasks. The code is available at the link https://github.com/microsoft/TimeCraft/tree/main/Diff-MN.
Keywords Time Series Generation Diffusion Models Mixture of Experts Neural Controlled Differential Equations Irregular Observations
@@ 1 Introduction
Time series generation (TSG) is essential across many real-world domains, including healthcare (Deng et al., 2025a, b), finance (Zhang et al., 2025a; Vasiliu et al., 2024; Zhang et al., 2024b), manufacturing and energy systems (Zhang et al., 2025b; Fuest et al., 2025; Zhang et al., 2025c). It supports data augmentation (Tang et al., 2025), privacy-preserving data sharing (Tian et al., 2024), and high-fidelity simulation for decision-making (Turowski et al., 2024). Despite substantial progress, most TSG methods assume regularly sampled TS, treating both training and generated data as discrete observations at fixed intervals. As shown in Fig. 1(a), existing approaches typically train on uniformly observed sequences and generate outputs at fixed resolutions (Mushunje et al., 2023; Li, 2023; Huang et al., 2025; Ge et al., 2025). While a few methods allow irregular training data, they still generate sequences at fixed intervals, as illustrated in Fig. 1(b) (Nikitin et al., 2024; Ramponi et al., 2018). This setting mismatches real-world scenarios, where observations are often irregular and sparse, yet downstream tasks require continuous and high-resolution TS. For example, in ICU monitoring, electronic health records are irregular and sparse, while applications such as patient trajectory simulation or risk forecasting require dense sequences (Tian et al., 2024; Li et al., 2025a). Consequently, continuous TSG in Fig. 1(c) from irregular observations remains an under-explored but critical challenge.
@@
Continuous-time modeling and irregular observations have been widely studied, leading to several approaches. Among them, Neural Ordinary Differential Equations (NODEs) and their extension, Neural Controlled Differential Equations (NCDE), are particularly popular. By combining continuous-time differential equations with learnable neural networks, NCDE-based methods effectively model temporal dynamics and handle irregular observations. Recent work has applied them to TSG (Lim et al., 2023; Naiman et al., 2024), typically following a two-stage paradigm: irregularly sampled TS are first transformed into regularly sampled ones using NCDE or their variants, and standard TSG models such as GANs (Jeon et al., 2022) and VAEs (Naiman et al., 2024) are then applied to generate data on regular time grids.
However, on the one hand, existing methods mainly use NCDE to process irregular inputs and then generate regularly sampled TS (Fig. 1(b)), with limited exploration of continuous TSG and evaluation (Fig. 1(c)). On the other hand, they directly adopt the standard NCDE without further improvement. Although standard NCDE are effective for modeling irregular TS, they rely on a single dynamics function, lack dynamics-focused optimization design, and cannot generalize learned dynamics to newly generated samples from a generative model. As a result, there remains substantial room to improve standard NCDE for irregular TS modeling and high-quality continuous-time generation.
In this study, we propose Diff-MN, a framework for continuous TSG with irregular observations. Our work differs from existing studies in the following aspects:
1.
We are the first to formally study continuous TSG and evaluate it under the generated high-frequency TS data from two perspectives: TS forecasting performance and analytical solutions of cubic polynomials.
2.
To enable effective continuous TSG, we improve standard NCDE in two key ways. (i) We replace the single dynamics function with a mixture-of-experts (MoE) dynamics function, where different experts capture distinct temporal patterns and are aggregated to model complex dynamics. The learned MoE weights explicitly encode temporal dynamics, providing a basis for using the diffusion model to parameterize them, thereby generalizing to newly generated samples. (ii) We propose a dynamics-focused, decoupled optimization strategy for NCDE. Specifically, we replace the State Initialization Network (SIN) and Readout Network (RN) with a pre-trained channel-wise autoencoder, which warm-starts training and allows the model to focus on optimizing the dynamics function.
3.
Based on (i) and (ii), the MoE weights learned by the trained NCDE encode temporal dynamics that are specific to the training samples. When new samples are generated, their temporal dynamics may differ, leading to a mismatch with these sample-specific weights. To address this, we propose to parameterize the MoE weights for each new sample. Specifically, we jointly model TS data and their corresponding MoE weights with a diffusion model to learn their joint distribution. As a result, for each newly generated sample, matched MoE weights that reflect the true dynamics are generated and substituted back into the pre-trained NCDE, enabling more accurate continuous time series generation.
Our contributions are as follows:
We are the first to formally study continuous TSG and evaluate it on generated high-frequency time series, filling an important research gap.
@@
For better continuous TSG, we first propose a decoupled Mixture-of-Experts Neural CDE (MoE-NCDE), which enhances standard NCDE with the MoE dynamic functions and dynamics-focused optimization design. This enables NCDE to better capture complex and dynamically changing temporal patterns from irregular observations.
We further propose Diff-MN, a continuous TSG framework that jointly models TS data and MoE-NCDE temporal dynamics parameters (MoE weights) using a diffusion model. This design parameterizes the MoE weights and allows sample-specific weights to be generated and substituted back into a pre-trained MoE-NCDE, enabling accurate continuous TSG for newly generated samples.
We conduct extensive experiments on four popular time-series forecasting datasets, four medical ECG datasets, and two synthetic datasets. Both quantitative and qualitative results demonstrate that Diff-MN consistently outperforms strong baselines on both irregular-to-regular and irregular-to-continuous generation tasks.
@@ 2 Related work
The most relevant methods are NCDE-based generators, e.g., KOVAE (Jeon et al., 2022) and GTGAN (Naiman et al., 2024), which build on standard NCDE and conduct irregular-to-regular generations. Some approaches also attempt to use attention-style mechanisms to capture the relationship between the observed sparse values and the dense time grids, learning the time dependencies and thereby enabling prediction at arbitrary time-steps (Shukla and Marlin, ; Yalavarthi et al., 2025). Unlike existing methods, we are the first to formally study the continuous TSG task and propose the Diff-MN framework, which introduces a mixture-of-experts dynamics function, a decoupled NCDE architectural design, and diffusion-based parameterization of MoE weights to generate sample-specific NCDE dynamics for each newly generated TS. Complete related work about TSG models for regular and irregular data is provided in the Appendix A.
@@ 3 Method
Overview. Our method is shown in Figure 2. We next detail the training and generation process of Diff-MN.
Training: First, we train the decoupled MoE NCDE on irregular observations, learning shared State Initialization Network (SIN), Readout Network (RN), and dynamic-function neural networks, along with sample-specific MoE weights that encode each instance’s temporal dynamics. These components are then used to impute missing values. Next, to parameterize the MoE weights for generalizing to new samples, we jointly train a diffusion model on the imputed data and their corresponding MoE weights. Consequently, the diffusion model generates new samples together with matched MoE weights that reflect their temporal dynamics, enabling better continuous TSG.
Generation: The newly generated samples and MoE weights from the diffusion model are fed back into pretrained MoE-NCDE for continuous TSG, producing high-frequency, longer, and more informative TS.
3.1 Preliminary
@@
Notation. Suppose we have observed multivariate TS. Each TS , for , is a continuous multivariate function defined over its time interval, where is the number of channels/variables. In practice, we observe each multivariate TS at discrete and irregular time points , where . The observed dataset is:
@@
Problem Definition. Given dataset , the goal is to learn a generative model over discrete and irregular TS for irregular or continuous generation, i.e., our goal is to learn a model capable of sampling new sequences . The generated dataset is denoted as where denotes arbitrary time step of -th sample.
Neural Controlled Differential Equation. Neural CDE (Kidger et al., 2020) extends Neural ODE (Chen et al., 2018) by modeling latent dynamics driven by a continuous control path. In continuous TSG, values generated between observations must align with surrounding temporal dynamics to avoid deviating from the true system distribution. Theoretically, we use NCDE instead of NODE because NCDE conditions the dynamics on irregular observations via a control path, enabling more accurate interpolation, whereas NODE may drift without observation guidance. The details of why NCDE is better in our task from both theoretical and experimental perspectives can be found in Appendix B.1.
The NCDE is defined as:
@@
where is a learnable dynamic-function neural network, and is a continuous interpolation (e.g., a natural cubic spline) of the discrete observations . is the initial latent state mapped from based on the network SIN.
Notably, although the control path is smooth due to spline interpolation, the latent trajectory can still be non-smooth because its evolution is governed by the nonlinear neural network and the mixture-of-experts, which together allow expressive, non-smooth, and highly diverse temporal patterns. The proof from both theoretical and experimental perspectives can be found in Appendix B.2.
Finally, CDESolver to solve NCDE is formalized as:
Its output is the latent state at arbitrary time points . Then, the network RN subsequently maps this hidden representation back to the original channel space to produce the predicted TS values.
3.2 Modeling diverse dynamics with dense MoE
Real-world TS exhibit diverse temporal patterns that a single dynamic function often struggles to capture. Using a Mixture of Experts (MoE), the complex temporal dynamics are first decomposed, allowing each expert to focus on the patterns it is best suited to learn, such as local or global trends. This makes the overall dynamics easier to learn while enhancing both expressiveness and generalization. By combining their outputs, the MoE-NCDE produces richer and more flexible temporal representations. This is strongly supported by Figure 6 in the experimental section.
Notably, we use a dense MoE rather than a sparse one, prioritizing stability over efficiency. Dense MoE fully engages all experts at each step, enabling stable training and better learning of complex temporal dynamics in our task, which involves limited training data and low-parameter models.
@@ 3.2 Modeling diverse dynamics with dense MoE
The MoE dynamics consist of a MLP (Multilayer Perceptron) based router and dynamic functions (each parameterized as a neural network), where . The router takes (potentially irregular) observations as input and outputs a routing score for each expert. Specifically, we first impute the irregular TS using cubic interpolation and then feed them into to obtain all experts’ scores for the -th sample. The router is implemented as a simple MLP. A softmax function is then applied over the scores to produce a dense routing distribution. Given , the final output is computed as a weighted combination of all expert functions:
Finally, substituting into the standard NCDE formulas (Eq. 2 and Eq. 3) yields the MoE-NCDE.
Notably, the learned MoE weights explicitly encode the temporal dynamics of each training sample. This property provides a foundation for parameterizing the MoE via a diffusion model. Consequently, for newly generated samples, the diffusion model can produce adapted MoE weights that capture their temporal dynamics, enabling customized and effective continuous TSG. A detailed description will be provided in Section 3.4.
3.3 Decoupled optimization design for MoE-dynamics-focused training
As shown in Figure 2(a), training a standard NCDE requires jointly optimizing the State Initialization Network (SIN), the Readout Network (RN), and the dynamic function network. Joint training can cause conflicts among modules and make the dynamic function network, which is the most critical component, difficult to learn effectively, especially when SIN and RN are poorly initialized.
To address this, we pretrain a channel-wise autoencoder on the irregular data and replace SIN and RN with the pretrained Channel-wise Encoder (CE) and Decoder (CD), keeping them frozen. This decoupled optimization (termed Decoupled Design) enables training to focus on the optimization of MoE dynamic functions while using CE and CD for a warm start, providing effective gradients and improving training stability. This is feasible because the pretrained CE and CD have the same input and output dimensions as SIN and RN, allowing direct substitution. The training of the CE and CD is shown in Appendix Algorithm 1 and overall MoE-dynamics-focused training in Appendix Algorithm 2.
3.4 Diffusion parameterized MoE-NCDE for continuous TSG
Building on the MoE-NCDE trained on a batch of samples , it can generate continuous time sequences by predicting values at any consistent with contextual time information. However, the learned MoE weights, which encode temporal dynamics, are specific to the training samples. New samples generated via a diffusion model have different temporal dynamics, rendering the original MoE weights incompatible. To address this, we propose parameterizing the MoE weights with a diffusion model, enabling newly generated samples to obtain adapted MoE weights that capture their temporal dynamics and allow customized, effective continuous TSG.
Specifically, the standard diffusion model requires regular data, but our training samples are irregular. We first use the pretrained MoE-NCDE to impute each sample into a regular sequence. The regularized samples, along with their corresponding MoE weights , are then fed into the diffusion model to learn the joint distribution of temporal patterns and MoE weights. This parameterizes the MoE weights, allowing new samples to receive adapted weights that describe their temporal dynamics. The training process is shown in the Appendix algorithm 3.
Finally, we input into the pre-trained MoE-NCDE and perform continuous TSG for each sample using its generated MoE weights and CDESolver. The process is shown in Appendix Algorithm 4.
@@ 4 Experiments
The experiment aims to answer the following questions: